Q)Let bx + ay =ab and ax+by=1 be two variable lines and a,b be parameters such that a 2 +b 2 =ab,then the locus of point of intersection of the lines is a)x 2 +y 2 +xy+1=0 b)x 2 +y 2 -xy+1=0 c)x 2 +y 2 +xy-1=0 d)x 2 +y 2 -xy-1=0

 
Q)Let bx + ay =ab and ax+by=1 be two     
variable lines and a,b be parameters such that a2+b2=ab,then the locus of point of intersection of the lines is
a)x2+y2+xy+1=0
b)x2+y2-xy+1=0
c)x2+y2+xy-1=0
d)x2+y2-xy-1=0 

Grade:11

1 Answers

UPINDERJEET SINGH
9 years ago
Let (p,q) be their point of intersection.
therefore:
bp+aq=ab
& ap+bq=1
multiply both,
abp^2+b^2pq+a^2pq+abq^2=ab
ab(p^2+q^2) + pq(a^2+b^2)=ab
but given (a^2+b^2)=ab
so,
ab(p^2+q^2+pq)=ab
thus,
p^2+q^2+pq-1=0
therefore locus is: x^2+y^2+xy-1=0 ie. c)